OpenStudy (anonymous):

these are hard to put on here but what is the solution to the system of equations |{ -3x-3y+2z=-7 <|{ z=1 |{ -2x-3y+z=-6

OpenStudy (anonymous):

first plug z = 1 into equns. 1 and 3 to give 2 equns. in x and y which u can solve by elimination

OpenStudy (anonymous):

-3x -3y +2 = -7 -2x -3y +1 = -6 subtractng the above eqautions : -x +1 = -1 x = 2 Plug x=2 in second equn: -2(2) -3y + 1 = -6 -3y = -6 -1 +4 = -3 y = 1 Solution is x=2, y=1 and z =1

OpenStudy (anonymous):

\[-3x-4y+2=-7\] \[-2x-3y+1=-6\] first is \[-3x-4y=-9\] second is \[-2x-3y=-7\] first is \[3x+4y=9\] second is \[-2x-3y=-7\] multiply first by 2 and second by 3 \[6x+8y=18\] \[-6x-9y=-21\] add tp get \[-y=-3\] or \[y=3\]

OpenStudy (anonymous):

oops maybe i am wrong sorry . let me check

OpenStudy (anonymous):

i may be wrong.

OpenStudy (anonymous):

jimmy rep is right and i am wrong. i will see why

OpenStudy (anonymous):

oh because i wrote -4y and it is -3y. my mistake. sorry